Show that the matrix [ 4 6 6 1 −1 3 2 −5 −2 ] is not diagonalizable


1
Expert's answer
2021-06-02T11:27:50-0400
"\begin{pmatrix}
 4& 6 & 6 \\
 1 & -1 & 3 \\
2 & -5 & -2 \\
\end{pmatrix}"

Start from forming a new matrix by subtracting "\lambda"  from the diagonal entries of the given matrix:


"\begin{pmatrix}
 4-\lambda & 6 & 6 \\
 1 & -1-\lambda & 3 \\
2 & -5 & -2-\lambda \\
\end{pmatrix}"

"\begin{vmatrix}
 4-\lambda & 6 & 6 \\
 1 & -1-\lambda & 3 \\
2 & -5 & -2-\lambda \\
\end{vmatrix}"

"=(4-\lambda)\begin{vmatrix}
 -1-\lambda & 3 \\
 -5 & -2-\lambda
\end{vmatrix}-6\begin{vmatrix}
 1 & 3 \\
 2 & -2-\lambda
\end{vmatrix}"

"+6\begin{vmatrix}
 1 & -1-\lambda \\
 2 & -5
\end{vmatrix}"


"=(4-\lambda)(1+\lambda)(2+\lambda()+15(4-\lambda)"

"+6(2+\lambda)+36-30+12(1+\lambda)"

"=8+12\lambda+4\lambda^2-2\lambda-3\lambda^2-\lambda^3+60-15\lambda"

"+12+6\lambda+6+12+12\lambda"

"=-\lambda^3+\lambda^2+13\lambda+98"

Solve the equation 


"-\lambda^3+\lambda^2+13\lambda+98=0"

Only approximate roots can be found.

The roots are 

"\lambda_1\approx5.9513233156912"


"\lambda_2\approx−2.4756616578456+3.21527996418223i"


"\lambda_3\approx−2.4756616578456-3.21527996418223i"


Since there are complex roots, the matrix is not diagonalizable.